Respond to the following question:

Suppose a country has a national debt of $5,000 billion, a GDP of $10,000 billion, and a budget deficit of $100 billion.

How much will its new national debt be?
Compute its debt-GDP ratio.
Suppose its GDP grows by 1% in the next year and the budget deficit is again $100 billion. Compute its new level of national debt and its new debt-GDP ratio.
Explain your answers for all of these questions (one well composed paragraph for each question).
1
Expert's answer
2015-03-20T12:19:24-0400
National debt of $5,000 billion, a GDP of $10,000 billion, and a budget deficit of $100 billion.
Its new national debt is 5,000 + 100 = $5,100 billion.
Its debt-GDP ratio is 5,100/10,000*100% = 51%, so the level of debt is affordable for the country (less then 60% of GDP).
If its GDP grows by 1% in the next year and the budget deficit is again $100 billion, the new level of national debt will be 5,100 + 100 = $5,200 and its new debt-GDP ratio will be 5,200/10,100*100% = 51.5%, so the level of debt is still affordable (less then 60%).